1. c) The radius of a silver bar is 30 cm. What force is required to stretch the bar by 12.7% of its length where Young's modulus is Y = 7.25 x 10¹0 Nm-².


1
Expert's answer
2021-12-06T09:46:23-0500

The strain is 12.7%, or 0.127, which can be expressed as:


"\epsilon=\frac{\sigma}{E}=\frac{F}{AE},\\\space\\
F=\epsilon AE=\epsilon·\pi r^2·E,\\\space\\
F=2.6\text{ MN}."